two cyclists start from the same point at the same time but went in separate directions. after 6 hours they were 126 miles apart. the rate of the second cyclist was 4 mph faster than the first. how fast was each one traveling
Answer by TimothyLamb(4379)

s = d / t
---
s = 126 / 6 = 21 mph
---
b = a + 4
a + b = 21
---
a + b = 21
a + a + 4 = 21
2a = 17
---
a = 8.5 mph
b = 12.5 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php